1/3

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

Der Werwolf La lupfantomo The Banshee (An Approach)

The Werewolf

tradukita de Bertilo Wennergren tradukita de Max Knight

tradukita de Alexander Gross

Ein Werwolf eines Nachts entwich von Weib und Kind und sich begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge mich!“ Jen lupfantom’ de famili’ meznokte iam kuris for al tombo de la profesor’. kaj petis pri konjugaci’.

One night, a banshee slunk away from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messing-

knauf und sprach zum Wolf, der seine

Pfoten geduldig kreuzte vor dem Toten:

Do diktis tiu dokta ul’ starante sur la kructabul’ dum lup’ senmova kiel splinto atentis vortojn de l’ mortinto.

The village teacher climbed up

straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was

ready.

�Der Werwolf“ – sprach der gute

Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s

nennt, den Werwolf, - damit hat’s ein

End!“ De “lupfantomo”, simpla rek-

ta, tra “lupfintomo”, plej perfek-

ta, plu “lupfontomo”, iom rara, ĝis “lupfuntomo”, ho,

koŝmara.

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

…  …  …  …

2/3

Dem Werwolf schmeichelten die

Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“ Ĉi participoj plaĉis multe al la fantomo, kaj rezulte ĝi petis: “Legu plu el PIV-o pri participoj en pasivo.”

The banshee marveled at the cases and writhed with pleasure, making fa-

ces, but said: “You did not add, so far, the plural to the singular!”

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts wus-

ste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

La viro ĉagreniĝis tamen: “Jam fermis mi per fin’ kaj

amen, ĉar netransiraj bestoj feblas kaj la pasiv’ neniam eblas.”

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

“A ‘Will-be-wolf?’ It’s just too long:

‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly

wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulula.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the te-

acher.

The Werewolf knew better–his sons still

slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

…  …  …  …

3/3

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en Esperanton de Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu la retejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .